# Indoor Location Fingerprinting Privacy: A Comprehensive Survey

### v1

@@ 6.1.2. Cryptography-based ILFPPM
Cryptography-based ILFPPM uses encryption to protect users’ indoor locations. Encryption is the most popular mechanism, probably because it is a standard solution for securely transmitting data. Fig. 8 illustrates how can Homomorphic encryption and SMPC are utilized in indoor location fingerprinting systems.
Table 1 reports all the studies that employ Cryptography for ILFPPM. As can be seen, authors in (Li et al., 2014) present the first indoor location privacy preservation work by encrypting measured RSS. It also introduces the attack models for data privacy and location privacy which have been followed by other studies as well. However, it was found insecure in (Yang and Järvinen, 2018). Most papers’ localization process is done through kNN algorithm as can be seen in the Table; however, Support Vector Machine (SVM) (Zhang et al., 2016) and fuzzy logic (Wang et al., 2015) are also applied for probabilistic location estimation. As can be seen different cryptography-based approaches including Paillier cryptosystem, Garbled Circuits (GC), and DGK algorithm are applied in ILFPPMs. Most papers employ RSS values but the application of cryptography for CSI data is also introduced in (Wang et al., 2015). The most common metrics applied to evaluate the proposed schemes are computation and communication overhead and localization error. Jarvinen et al. propose PILOT (Järvinen et al., 2019), which was the first efficient solution for ILFPPM. PILOT employs well-known Secure Two-Party Computation (STPC) techniques and outsources the most involving computations to two semi-trusted servers to achieve practical performance. Zhang et al. (Zhang et al., 2016) shows a PPIL based on SVM and Paillier encryption, but (Yang and Järvinen, 2018, 2018) shows that it suffers from similar weaknesses. They discuss the state-of-the-art active attacks (theoretically) and server data privacy attacks. As a solution, they represent GC theoretically. In another work, Nieminen and Jarvinen (Nieminen and Järvinen, 2021) proposed a PPIL scheme in the user-server setting using Paillier encryption and GCs.
@@
Cryptography methods operate without relying on TTP or outsourcing but they involve relatively expensive operations. Their scheme requires computationally heavy pre-computations in the setup phase for each localization query. These computations, especially on user devices, demand resources, posing a significant concern for devices with limited battery life.
6.2. Anonymization
@@
Here, we briefly explain the anonymization techniques and then discuss the studies on anonymization-based ILFPPM.
@@
Figure 9. Examples of $3$-anonymity for location and trajectory privacy protection
6.2.1. Anonymization techniques
@@
Given a dataset $\mathcal{D}$, a quasi-identifier $\mathcal{QI}$, and a parameter $k$, $\mathcal{D}$ satisfies $k$-anonymity if and only if, for each combination of quasi-identifier values in $\mathcal{QI}$, there exist at least $k$ records in $\mathcal{D}$ with identical quasi-identifier values.
Definition 6.4 ($\ell$-diversity (Machanavajjhala et al., 2006)).
@@
Given a dataset $\mathcal{D}$, a quasi-identifier $\mathcal{QI}$, and a parameter $\ell$, $\mathcal{D}$ satisfies $\ell$-diversity if, for each combination of quasi-identifier values in $\mathcal{QI}$, there are at least $\ell$ ”well-represented” sensitive attribute values in the corresponding group of records.
Definition 6.5 ($t$-closeness (Li et al., 2007)).
@@
Given a dataset $\mathcal{D}$, a quasi-identifier $\mathcal{QI}$, and a parameter $t$, $\mathcal{D}$ satisfies $t$-closeness if, for each combination of quasi-identifier values in $\mathcal{QI}$, the distribution of sensitive attribute values in the corresponding group of records is ”close” to the overall distribution of sensitive attribute values in the entire dataset.
Table 2. Studies on anonymization-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Adversary Model Attack Model (Kim et al., 2012) 2012 - $k$-anonymity Hierarchical graph Simulated Area of ASR No. of cells - Loc. Priv. Attack Inference Attack (Zhu and Zhai, 2014) 2014 kNN, SVM $k$-anonymity cloaking obfuscation Simulated Success Rate Area of ASR - Loc. Priv. Attack (Konstantinidis et al., 2015) 2015 - $k$-anonymity Bloom Filter Simulated CSUCY KIOSUCY Crawdad Energy Consumption No. of Messages Run Time Fully malicious Loc. & Data Priv. Attack Linking Attack (Kim and Li, 2016) 2016 kNN $k$-anonymity $\ell$-diversity Hierarchical graph Simulated Success Rate Area of ASR Response Time Fully trusted Loc. Priv. Attack Protocol-based Attack Context Linking Attack (Alikhani et al., 2018) 2018 NN Hilbert Curve $k$-anonymity Simulated Coms. Overhead Loc. Error - - (Sazdar et al., 2020b) 2020 kNN $k$-anonymity Randomization Permutation Simulated CRI Coms. Overhead Loc. Error Fully malicious Loc. Priv. Attack (Zhao et al., 2020) 2020 kNN $k$-anonymity Simulated Success Rate CR Area Coms. Overhead Loc. Error - Loc. Priv. Attack Context Linking Attack (Sazdar et al., 2020a) 2021 kNN $k$-anonymity Hilbert Curve Bloom Filter Simulated CRI CSUCY KIOSUCY PosData Loc. Error Semi-trusted Loc. priv. Attack (Fathalizadeh et al., 2022) 2022 - $k$-anonymity $\ell$-diversity $t$-closeness ($\alpha,k$)-anonymity $\delta$-presence Simulated CRI UJIIndoorLoc Spatial Loss RSS Loss NCP Semi-trusted Loc. & Data Priv. Attack Semantic Attack Data Linking Attack Probability-based Attack
6.2.2. Anonymization-based ILFPPM
Anonymization techniques can be employed in IPSs to protect the privacy of individuals while still enabling the provision of LBS. As illustrated in Fig. 9, these techniques can be applied to both location and trajectory data. The application of anonymization techniques for indoor localization was first introduced by authors in (Kim et al., 2012) using graph hierarchy, where nearby zones were mixed based on different algorithms. Due to $k$-anonymity’s vulnerability to identification and re-identification attacks, the proposed methods often combine with various approaches, such as Hilbert Curve, Bloom Filter, $\ell$-diversity, etc.
Table 2 reports the methods proposed for ILFPPMs based on anonymization, which indicates that the application of $k$-anonymity was first introduced by Hierarchical Graphs, followed by cloaking and obfuscation methods. The most widely used metric for analyzing these methods is the Success Rate to assess how often $k$-anonymity is satisfied. $k$-anonymity is also used for protecting the trajectory of the users inside the buildings (Sazdar et al., 2020b). The use of Hilbert Curve and Bloom Filter, both separately and in combination, is integrated with anonymization methods to enhance location privacy protection. In the latest work (Fathalizadeh et al., 2022) five methods are introduced, $k$-anonymity, $\ell$-diversity, $t$-closeness, ($\alpha,k$)-anonymity, $\delta$-presence, to defend against five major attacks.
6.3. Differential Privacy
@@
Here, we provide a brief explanation of the Differential Privacy (DP) mechanism and various DP-based solutions and subsequently discuss the DP-based ILFPPM.
6.3.1. DP-based techniques
Definition 6.6 ($\epsilon$-differential privacy (Dwork et al., 2006)).
[…]
(2) $\frac{\Pr[\mathcal{A}(\mathcal{D})\in S]}{\Pr[\mathcal{A}(\mathcal{D}^{\prime})\in S]}\leq e^{\epsilon},$
where, the probability is computed over the randomness used by the algorithm.
Figure 10. CDP vs LDP in IPS
Definition 6.7 (Geo-indistinguishability (Andrés et al., 2013)).
@@ 6.3.1. DP-based techniques
When releasing aggregated information about a group of individuals, DP is an effective tool. However, it may not be suitable for scenarios involving a single individual, such as location privacy. DP prevents any change in an individual’s location from affecting the published output, making it challenging to convey meaningful information to the service provider. To address these limitations, Geo-Indistinguishability was introduced. Geo-Indistinguishability adds random noise to a user’s true location to prevent an adversary with background knowledge from accurately inferring the user’s location. This approach utilizes 2D Euclidean distance, suitable for localization purposes, whereas previous methods primarily focused on Hamming distance. Formally, for a randomized mechanism $\mathcal{K}$ that reports an obfuscated location $z$ instead of the user’s real location $x$, $\epsilon$-Geo-Indistinguishability is satisfied if, for any location ${x}^{\prime}$:
(3) $\mathcal{K}(x)(z)\leq e^{\epsilon d(x,{x}^{\prime})}\cdot\mathcal{K}({x}^{\prime})(z),$
@@
where, $d$ represents the Euclidean distance between $x$ and ${x}^{\prime}$. Geo-Indistinguishability ensures that the probability of reporting point $z$ instead of other points $x$ and ${x}^{\prime}$ differs by at most $e^{-\epsilon d(x,{x}^{\prime})}$. This is achieved using a noise function, causing the probability of reporting a point around $z$ to decrease exponentially as the distance from the actual location $x$ increases. In a 1D linear space, this is accomplished using the Laplace distribution with the probability distribution function (PDF) $f_{X}(x)=\frac{\epsilon}{2}e^{-\epsilon\left|x-\mu\right|}$, where $\mu$ and $\epsilon$ are parameters in the Laplace distribution.
Table 3. Studies on different DP-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Adversary Model Attack Model (Li et al., 2016) 2016 kNN CDP Paillier Private Coms. Overhead Loc. Error Semi-trusted Loc. Priv. Attack (Sadhu et al., 2017) 2017 - LDP Two-step classifier Multi Building Rutgers Uni Loc. Error - - (Zhu et al., 2017) 2017 - LDP Delaunay triangulation Range generalization Simulated Loc. Error - - (Kim et al., 2018) 2018 - LDP Randomized Response Simulated Sangmyung Uni Error Rate Fully malicious Loc. Priv. Attack Probability-based Attack (Wang et al., 2018) 2018 - CDP AP fuzzification Finger Clustering Finger Permutation PosData Loc. Error Distance Error - Loc. & Data Priv. Attack Probability-based Attack (Zhao et al., 2018) 2018 - LDP Segmenting Data $k$-anonymity Private Coms. Overhead Loc. Error Semi-trusted Loc. Priv. Attack Context linking Attack Probability-based Attack Protocol-based Attack (Kim and Jang, 2019) 2019 - LDP Randomized Encoding Simulated Beijing Taxi Trajectory Loc. Error - - (Navidan et al., 2022) 2022 - LDP Local Hashing Unary Encoding Histogram Encoding Random response RAPPOR JUIndoorLoc CRI Loc. Error - Loc. & Data Priv. Attack (Zhang et al., 2022) 2022 FSELM LDP Edge–cloud collaboration Private Coms. Overhead Loc. Error Run Time Fully malicious Loc. Priv. Attack Probability-based Attack Inference Attack (Min et al., 2022) 2022 - GeoInd Laplace Mechanism Simulated Coms. Overhead Loc. Privacy QoS Loss Fully malicious Loc. Priv. Attack Probability-based Attack Inference Attack (Fathalizadeh et al., 2023) 2023 - GeoInd Laplace Mechanism Distance calculation RSS Generation Simulated JUIndoorLoc CRI Loc. Error Probabilistic QoS Loss Fully malicious Loc. Priv. Attack Probability-based Attack Context Linking Attack (Min et al., 2023) 2023 - GeoInd Reinforcement learning Simulated Coms. Overhead Loc. Privacy QoS Loss Fully malicious Inference Attack Semantic Attack Probability-based Attack (Hemkumar, 2024) 2024 - LDP GeoInd Geolife Gowalla QoS Loss Fully malicious Inference Attack Loc. Priv. Attack Probability-based Attack
6.3.2. DP-based ILFPPM

### v2

@@ 5.3.2 Comparative analysis of Cryptographic techniques
Cryptography-based ILFPPM uses encryption to protect users’ indoor locations. Encryption is the most popular mechanism, probably because it is a standard solution for securely transmitting data. Fig. 8 illustrates how Homomorphic encryption and SMPC are utilized in ILF systems.
Table 3 reports all the studies that employ Cryptography for ILFPPM. Authors in [44] presented the first indoor location privacy preservation work by encrypting measured RSS. It also introduces the attack models for data privacy and location privacy mentioned in section 4.2.1, which other studies have followed. However, the proposed method was found insecure in [45]. Additionally, Järvinen et al. introduced PILOT [75], the first efficient solution for privacy-preserving indoor localization using STPC, where the computational load is outsourced to two semi-trusted servers to achieve practical performance.
Most papers rely on the k-Nearest Neighbors (kNN) algorithm for the localization process, as highlighted in the Table; however, alternative methods have also been explored. For instance, Support Vector Machine (SVM) [72] has been employed for enhanced classification accuracy in localization tasks, while fuzzy logic [35] has been applied to handle uncertainty and imprecision in probabilistic location estimation. Additionally, Maximum Likelihood Estimation (MLE) [80] has been utilized to improve the accuracy of location predictions by maximizing the probability of the observed data. These approaches provide more flexibility in modeling complex environments, particularly in scenarios where the signal measurements exhibit significant variability or noise. For the data applied in simulations and experiments, most papers primarily use RSS values for localization; however, the integration of cryptographic techniques with CSI data has also been explored, as demonstrated in [35]. Commonly used metrics to assess the effectiveness of these proposed schemes include the computation and communication overhead, as well as the localization error. These metrics provide a comprehensive evaluation of both the security and performance trade-offs, ensuring that the cryptographic enhancements do not excessively degrade system efficiency while maintaining accurate location estimates.
The use of cryptographic techniques in privacy-preserving indoor location data has been explored in various ways across the surveyed studies, with most leveraging the Paillier cryptosystem, Garbled Circuits (GC), and the DGK algorithm. Paillier encryption is favored for its homomorphic properties, enabling encrypted data computations without decryption, which is commonly employed in studies such as [44], [35], and [77] to maintain location privacy with minimal computational overhead. However, its performance trade-offs, particularly in real-time localization systems, include latency and communication overhead, as noted by [78] and [64]. Similarly, Zhang et al. [72] present a privacy-preserving indoor localization (PPIL) system based on SVM and Paillier encryption, although it faces limitations due to reliance on semi-trusted servers, as discussed in [45]. In contrast, Garbled Circuits offer more robust security features and are typically applied in scenarios involving fully malicious adversaries. Studies such as [45] and [73] combine Garbled Circuits with Paillier encryption to strengthen privacy guarantees, especially when both location and data privacy are at risk. However, the higher computational costs of GC limit its suitability for low-latency applications, as noted by [72] and [79]. Nieminen and Järvinen [64] further explored the combination of Paillier encryption and Garbled Circuits in a PPIL scheme for the user-server setting.
While cryptographic methods such as Paillier and Garbled Circuits provide varying levels of privacy and efficiency, the choice of technique depends heavily on the adversary model, the dataset size, and the performance requirements of the application. A key observation across all studies is the diversity of datasets used, ranging from simulated environments to public datasets like JUIndoorLoc, Colombia University, and Tampere University. Studies based on real-world datasets, such as [73], [63], [78], and [81], provide more practical insights into the performance of these cryptographic methods in realistic settings. However, there is a need for more extensive testing on public datasets to standardize performance comparisons.
@@
Overall, cryptographic methods operate without relying on a TTP, but they often involve relatively expensive operations. While these methods can be adapted to both client-server and outsourced settings such as the outsourcing-to-two-servers model discussed in [82], they typically require computationally heavy pre-computations in the setup phase for each localization query. These computations, especially on user devices, need significant resources, which can be a concern for devices with limited battery life.
5.3.3 Discussion on employing Cryptography-based methods
To apply cryptographic techniques in ILF systems, the following guidelines will help achieve robust privacy protection while maintaining operational effectiveness:
Balancing privacy and localization accuracy: Methods like homomorphic encryption allow encrypted data processing without decryption but introduce significant computation delays. In scenarios where both precision and speed are critical in localization, SMPC can offer a more balanced approach. SMPC enables joint computations on private data while maintaining strong privacy guarantees and typically improves system responsiveness compared to fully homomorphic encryption. However, SMPC can introduce communication overhead, particularly in larger systems. Thus, it is important to evaluate the trade-off between acceptable precision loss, privacy protection, and the system’s ability to handle the computational and communication demands.
Adaptability to adversarial models: When dealing with fully malicious adversaries, more robust cryptographic methods like Paillier encryption combined with advanced privacy protocols are needed, as used in secure multi-party computations. For less severe adversarial models, lighter cryptographic techniques such as homomorphic encryption or secret sharing schemes can be utilized to reduce computational overhead while maintaining an acceptable level of privacy and security. These methods provide a balance between efficiency and protection, depending on the severity of the adversarial model.
Encryption overhead and scalability: Cryptographic methods introduce varying levels of computational and communication overhead, which may impact scalability in large systems. For instance, homomorphic encryption methods are more resource-intensive but offer higher privacy. Simpler techniques like secret sharing or lightweight secure multi-party computations are computationally lighter but may offer less robust protection. It is important to choose encryption schemes that align with the privacy requirements and scalability needs of the desired deployment.
5.4 Anonymization in ILFPPM
@@
Here, we briefly explain the anonymization techniques and then discuss the studies on anonymization-based ILFPPM. Finally, we provide guidelines on employing anonymization in ILFPPM.
@@
Figure 9: Examples of 33-anonymity for location and trajectory privacy protection
5.4.1 Anonymization techniques
@@
$k$-anonymity [83]: Given a dataset $\mathcal{D}$, a quasi-identifier $\mathcal{QI}$, and a parameter $k$, $\mathcal{D}$ satisfies $k$-anonymity if and only if, for each combination of quasi-identifier values in $\mathcal{QI}$, there exist at least $k$ records in $\mathcal{D}$ with identical quasi-identifier values.
@@
$\ell$-diversity [84]: Given a dataset $\mathcal{D}$, a quasi-identifier $\mathcal{QI}$, and a parameter $\ell$, $\mathcal{D}$ satisfies $\ell$-diversity if, for each combination of quasi-identifier values in $\mathcal{QI}$, there are at least $\ell$ ”well-represented” sensitive attribute values in the corresponding group of records.
@@
$t$-closeness [85]: Given a dataset $\mathcal{D}$, a quasi-identifier $\mathcal{QI}$, and a parameter $t$, $\mathcal{D}$ satisfies $t$-closeness if, for each combination of quasi-identifier values in $\mathcal{QI}$, the distribution of sensitive attribute values in the corresponding group of records is ”close” to the overall distribution of sensitive attribute values in the entire dataset.
Table 4: Studies on anonymization-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Adversary Model Attack Model
[86] 2012 -
$k$-anonymity
Hierarchical graph
Simulated
Area of ASR
No. of cells
Loc. Priv. Attack
Inference Attack
[87] 2014 kNN, SVM
$k$-anonymity
cloaking
obfuscation
Simulated
Success Rate
Area of ASR
- Loc. Priv. Attack
[55] 2015 -
$k$-anonymity
Bloom Filter
Simulated
CSUCY
KIOSUCY
Crawdad
Energy Consumption
No. of Messages
Run Time
Fully malicious
Loc. & Data Priv. Attack
Linking Attack
[88] 2016 kNN
$k$-anonymity
$\ell$-diversity
Hierarchical graph
Simulated
Success Rate
Area of ASR
Response Time
Fully trusted
Loc. Priv. Attack
Protocol-based Attack
Context Linking Attack
[89] 2018 NN
Hilbert Curve
$k$-anonymity
Simulated
Coms. Overhead
Loc. Error
[90] 2020 kNN
$k$-anonymity
Randomization
Permutation
Simulated
CRI
Coms. Overhead
Loc. Error
Fully malicious
Loc. Priv. Attack
[91] 2020 kNN $k$-anonymity Simulated
Success Rate
CR Area
Coms. Overhead
Loc. Error
Loc. Priv. Attack
Context Linking Attack
[92] 2021 kNN
$k$-anonymity
Hilbert Curve
Bloom Filter
Simulated
CRI
CSUCY
KIOSUCY
PosData
Loc. Error Semi-trusted Loc. priv. Attack
[49] 2022 -
$k$-anonymity
$\ell$-diversity
$t$-closeness
($\alpha,k$)-anonymity
$\delta$-presence
Simulated
CRI
UJIIndoorLoc
Spatial Loss
RSS Loss
NCP
Semi-trusted
Loc. & Data Priv. Attack
Semantic Attack
Data Linking Attack
Probability-based Attack
5.4.2 Comparative analysis of anonymization techniques
Anonymization techniques in IPSs can protect individuals’ privacy by masking identifiable information, while still allowing LBS to function. As depicted in Fig. 9, these methods can be used for both location data (such as geographic coordinates) and trajectory data (movement over time). By anonymizing this data, IPSs lower the chances of re-identification, ensuring that sensitive information remains secure without compromising system performance.
Table 4 reports the methods proposed for ILFPPMs based on anonymization. One of the earliest works by [86] and [87] utilized $k$-anonymity combined with hierarchical graphs and cloaking mechanisms. These methods ensure that each individual is indistinguishable from at least $k$ other individuals within the dataset, mitigating the risk of re-identification. In these studies, performance is typically measured using metrics like Area of Anonymity Set Radius (ASR) and the number of cells, which indicate how effectively the location data is obfuscated. However, these methods may struggle to maintain high privacy guarantees against sophisticated adversaries, as they primarily target location privacy attacks and inference attacks, without addressing more advanced attack models like linking or semantic attacks.
In later studies, such as [55], $k$-anonymity is combined with Bloom filters to improve computational efficiency and scalability. This approach was evaluated using a variety of datasets, such as CSUCY and Crawdad, allowing the method to demonstrate its effectiveness in diverse environments. These methods showed improvements in terms of energy consumption, number of messages, and run time, which are critical in real-time indoor localization systems. However, the fully malicious adversary model adopted in this study highlights the vulnerability of these methods to more sophisticated attacks like linking and data privacy attacks. Therefore, while Bloom filters enhance efficiency, they require additional mechanisms to strengthen privacy guarantees in more hostile environments.
As privacy concerns grew, more advanced models like $\ell$-diversity and $t$-closeness emerged to address some of the inherent weaknesses of $k$-anonymity, particularly concerning attribute disclosure risks. For instance, [Kim 2016] integrated $\ell$-diversity into their $k$-anonymity framework, ensuring the diversity of sensitive attributes within each anonymity group is preserved. This model performed well against context-linking and protocol-based attacks while maintaining a high success rate and low response time. However, these methods typically assume a fully trusted adversary model, which may limit their effectiveness in real-world scenarios where such trust is unrealistic.
Recent works, such as [49] and [92], have extended anonymization techniques by combining $k$-anonymity with spatial methods like Hilbert curves and randomization techniques. These studies highlight a growing trend toward improving the resilience of anonymization methods in dynamic environments with semi-trusted or fully malicious adversaries. The use of randomization and permutation in these works provides additional layers of security against location privacy attacks, though these methods often introduce higher communication overhead and location errors. Furthermore, [49] adopts a comprehensive approach by incorporating multiple privacy models, such as $t$-closeness and $\delta$-presence, enabling protection against a wider range of attacks, including semantic and data linking attacks.
5.4.3 Discussion on employing Anonymization-based methods
To implement anonymization techniques in ILF, the following recommendations will help ensure both privacy protection and operational efficiency:
Selecting the right anonymity model: Depending on the specific privacy needs and the strength of the adversary model, a range of anonymization techniques can be utilized in IPSs. One commonly used approach is $k$-anonymity, which ensures that an individual’s data cannot be distinguished from at least k1k-1 others. However, $k$-anonymity may fall short against more sophisticated adversaries capable of performing inference attacks or cross-referencing external datasets. To address these vulnerabilities, enhanced models such as $\ell$-diversity or $t$-closeness can be implemented. $\ell$-diversity extends $k$-anonymity by ensuring that sensitive attributes within a group are diverse enough to prevent inference attacks, while $t$-closeness further strengthens protection by maintaining the distribution of sensitive data close to its overall population distribution, thereby mitigating risks of linking attacks and preventing adversaries from inferring sensitive information based on statistical imbalances. These extended techniques offer more robust privacy safeguards, especially in scenarios involving powerful adversaries or complex datasets.
Efficiency considerations: The additional overhead brought by anonymization techniques, including the computational requirements of obfuscating data and the increased energy consumption, must be thoroughly assessed when designing IPSs. These costs can significantly impact the performance of the system, particularly in terms of processing speed, battery life in mobile devices, and the efficiency of data transmission. Some methods, such as Hilbert Curve anonymization, are specifically designed to minimize these overheads. By mapping spatial data onto a one-dimensional curve, Hilbert Curve anonymization reduces communication costs and decreases localization errors, making it highly effective for real-time applications that require both privacy protection and low-latency performance. This balance between privacy and system efficiency makes such methods ideal for deployment in resource-constrained environments, such as mobile devices or large-scale, real-time IPS networks. However, selecting the right anonymization technique requires careful consideration of trade-offs between privacy, computational overhead, and system responsiveness.
Trade-offs between privacy and utility: The utility of anonymized data can be significantly diminished due to data suppression or generalization techniques employed to protect privacy. These modifications often degrade the accuracy of the data, making it less useful for applications like IPS or LBS. To maintain an effective balance between privacy and functionality, it is essential to evaluate metrics such as location error, which measures the deviation from actual position, and run time, which affects system performance. By analyzing these metrics, the optimal trade-off between safeguarding privacy and retaining the usability of the data can be determined. Techniques such as Hierarchical Graphs [86] [88] and Hilbert Curves [89] [92] are particularly effective in achieving this balance. Hierarchical Graphs allow for varying levels of data abstraction, which can be adjusted based on privacy needs, while still preserving important spatial relationships. Hilbert Curves, on the other hand, provide a space-filling curve that helps reduce location error while minimizing the impact of noise or suppression. Both methods offer superior trade-offs, ensuring higher location accuracy while simultaneously enhancing privacy, making them well-suited for privacy-preserving IPS where data utility is crucial.
5.5 Differential Privacy in ILFPPM
@@
Here, we provide a brief explanation of the Differential Privacy (DP) mechanism and various DP-based solutions and subsequently discuss the DP-based ILFPPM. Finally, the guidelines on using DP for ILF are suggested.
5.5.1 DP-based techniques
$\epsilon$-differential privacy [93]: Consider a positive real number $\epsilon$ and a randomized algorithm denoted as $\mathcal{A}$, which takes a dataset as input, representing the actions of a trusted party holding the data. Let $\textit{im}(\mathcal{A})$ represent the set of possible outputs of the algorithm. The algorithm $\mathcal{A}$ is considered to provide $\epsilon$-differential privacy if, for all pairs of datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ that differ in a single element (i.e., the data of one person), and for all subsets $S$ of $\textit{im}(\mathcal{A})$, the following inequality holds:
$\frac{\Pr[\mathcal{A}(\mathcal{D})\in S]}{\Pr[\mathcal{A}(\mathcal{D}^{\prime})\in S]}\leq e^{\epsilon},$ (2)
where the probability is computed over the randomness used by the algorithm.
Figure 10: Centralized Differential Privacy vs. Local Differential Privacy in Indoor Positioning Systems
@@ 5.5.1 DP-based techniques
Geo-indistinguishability [94]: When releasing aggregated information about a group of individuals, DP is an effective tool. However, it may not be suitable for scenarios involving a single individual, such as location privacy. DP prevents any change in an individual’s location from affecting the published output, making it challenging to convey meaningful information to the service provider. To address these limitations, Geo-Indistinguishability was introduced. Geo-Indistinguishability adds random noise to a user’s true location to prevent an adversary with background knowledge from accurately inferring the user’s location. This approach utilizes 2D Euclidean distance, suitable for localization purposes, whereas previous methods primarily focused on Hamming distance. Formally, for a randomized mechanism $\mathcal{K}$ that reports an obfuscated location $z$ instead of the user’s real location $x$, $\epsilon$-Geo-Indistinguishability is satisfied if, for any location ${x}^{\prime}$:
$\mathcal{K}(x)(z)\leq e^{\epsilon d(x,{x}^{\prime})}\cdot\mathcal{K}({x}^{\prime})(z),$ (3)
@@
where $d$ represents the Euclidean distance between $x$ and ${x}^{\prime}$. Geo-Indistinguishability ensures that the probability of reporting point $z$ instead of other points $x$ and ${x}^{\prime}$ differs by at most $e^{-\epsilon d(x,{x}^{\prime})}$. This is achieved using a noise function, causing the probability of reporting a point around $z$ to decrease exponentially as the distance from the actual location $x$ increases. In a 1D linear space, this is accomplished using the Laplace distribution with the probability distribution function (PDF) $f_{X}(x)=\frac{\epsilon}{2}e^{-\epsilon\left|x-\mu\right|}$, where $\mu$ and $\epsilon$ are parameters in the Laplace distribution.
Table 5: Studies on different DP-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Adversary Model Attack Model
[95] 2016 kNN
CDP
Paillier
Private
Coms. Overhead
Loc. Error
Semi-trusted Loc. Priv. Attack
[61] 2017 -
LDP
Two-step classifier
Multi Building
Rutgers Uni
Loc. Error - -
[96] 2017 -
LDP
Delaunay triangulation
Range generalization
Simulated Loc. Error - -
[60] 2018 -
LDP
Randomized Response
Simulated
Sangmyung Uni
Error Rate Fully malicious
Loc. Priv. Attack
Probability-based Attack
[57] 2018 -
CDP
AP fuzzification
Finger Clustering
Finger Permutation
PosData
Loc. Error
Distance Error
Loc. & Data Priv. Attack
Probability-based Attack
[97] 2018 -
LDP
Segmenting Data
$k$-anonymity
Private
Coms. Overhead
Loc. Error
Semi-trusted
Loc. Priv. Attack
Context linking Attack
Probability-based Attack
Protocol-based Attack
[98] 2019 -
LDP
Randomized Encoding
Simulated
Beijing Taxi
Trajectory
Loc. Error - -
[99] 2022 -
LDP
Local Hashing
Unary Encoding
Histogram Encoding
Random response
RAPPOR
JUIndoorLoc
CRI
Loc. Error - Loc. & Data Priv. Attack
[100] 2022 FSELM
LDP
Edge–cloud
collaboration
Private
Coms. Overhead
Loc. Error
Run Time
Fully malicious
Loc. Priv. Attack
Probability-based Attack
Inference Attack
[101] 2022 -
GeoInd
Laplace Mechanism
Simulated
Coms. Overhead
Loc. Privacy
QoS Loss
Fully malicious
Loc. Priv. Attack
Probability-based Attack
Inference Attack
[102] 2023 -
GeoInd
Laplace Mechanism
Distance calculation
RSS Generation
Simulated
JUIndoorLoc
CRI
Loc. Error
Probabilistic
QoS Loss
Fully malicious
Loc. Priv. Attack
Probability-based Attack
Context Linking Attack
[103] 2023 -
GeoInd
Reinforcement learning
Simulated
Coms. Overhead
Loc. Privacy
QoS Loss
Fully malicious
Inference Attack
Semantic Attack
Probability-based Attack
[41] 2024 -
LDP
GeoInd
Geolife
Gowalla
QoS Loss Fully malicious
Inference Attack
Loc. Priv. Attack
Probability-based Attack
5.5.2 Comparative analysis of DP-based techniques